### v1

Abstract
[…]
keywords:
decoupling fusion, multi-modal fusion, IVF-ET pregnancy prediction.1 Introduction
Recent study shows that up to 12-15 of couples are diagnosed as infertility [Hornstein(2016)], and in vitro fertilization embryo transfer (IVF-ET) is one of the most effective technologies to treat infertility. As shown in Fig. 1, during the IVF-ET process, medical laboratory technicians obtain multiple oocytes by stimulating mother’s uterus with ovulation and produce multiple zygotes in a laboratory environment [Hanevik and Hessen(2022)]. After 3-5 days culture, laboratory technicians select the optimal embryos based on visual evaluation of embryo morphology and transfer it back to mother’s uterus for further development. Thus, it is a crucial step to select high-quality embryos that would lead to promising pregnancy results of IVF-ET.
In clinical practice, the pregnancy success rate of IVF-ET is 30-40 only [Gleicher et al.(2019)Gleicher, Kushnir, and Barad]. One reason is that the optimal embryos may not be survival after the transfer, since human evaluation of embryo morphology is highly subjective and with low consistency. Moreover, embryo morphology is not always relevant to embryos’ true development vitality. Numerous studies have shown that embryos with good morphology didn’t survive while ones with poor morphology did [del Carmen Nogales et al.(2021)del Carmen Nogales, Cruz, de Frutos, Martínez, Gaytán, Ariza, Bronet, and Garcia-Velasco]. In fact, not only the embryo morphology affects the pregnancy success rate, but also the fertility indicators of parents, such as parents’ age, endometrial thickness, sperm quality, and so on.
In the area of computer assisted IVF-ET, existing researches mainly focus on the morphological grading of embryos. As shown in Fig. 1, we can obtain microscope images of embryonic development between step 3 and step 4 of IVF-ET. In order to perform the embryos morphological grading task, [Wu et al.(2021)Wu, Yan, Li, Li, Wang, Chang, Yu, Jin, Ma, Luo, et al.] and [Liu et al.(2023a)Liu, Li, Dai, Shan, Zhang, Zhuang, Lee, Wong, Yue, Huang, et al.] apply convolutional neural networks (CNN) based on static embryo images; [Wang et al.(2024)Wang, Wang, Gao, Chen, Gao, Ma, Jiang, Yang, Feng, Zhang, et al.] and [Lukyanenko et al.(2021)Lukyanenko, Jang, Wei, Struyven, Kim, Leahy, Yang, Rush, Ben-Yosef, Needleman, et al.] apply transformer and two-stream neural network based on time-lapse microscopy (TLM) images, respectively. In addition, [Cheng et al.(2024)Cheng, Chen, Hu, Meng, Liu, Chen, Wu, and Ying] fuse multi-focal images to predict outcome of blastocyst. The performance of these methods outperform laboratory technicians, because embryos grading is completely based on morphological information and the salient morphological characteristics among different grades are easily distinguishable for machines. However, morphological grading is indirect and less relevant to the pregnancy outcome of IVF-ET as discussed in the previous paragraph. Therefore, more researches tend to predict the pregnancy outcome directly.
Recent AI-based assessment models achieve promising success in direct pregnancy predction. [Dehghan et al.(2024)Dehghan, Rabiei, Choobineh, Maghooli, Nazari, and Vahidi-Asl] apply traditional machine learning methods with fertility table indicators obtained before step 5 for pregnancy prediction. In addition, static images of the fifth day’s embryos and TLM images are also adopted to predict the pregnancy outcome, respectively [Kim et al.(2024)Kim, Ko, Kang, Choi, Park, Chung, Kim, Kim, and Lee, Berntsen et al.(2022)Berntsen, Rimestad, Lassen, Tran, and Kragh]. What’s more, [Liu et al.(2022)Liu, Zhang, Gu, Dai, Shan, Song, Li, Chen, Lin, and Sun](MMBE) fuse the fifth day’s static embryo image and fertility table indicators to achieve better pregnancy prediction performance. The major limitation of the existing image-based methods is that they are only applicable to the fifth day’s embryo transfer. However, in reality many reproductive centers carry out the third day’s embryo to transfer. Although the embryo of the third day is less developed than that of the fifth day, the embryo images of the third day can still provide clinically significant information for pregnancy prediction [Neblett et al.(2021)Neblett, Kim, Jones, Baumgarten, Coddington, Zhao, and Shenoy]. On the other hand, due to some technical constraints in reality multi-modal fusion method [Liu et al.(2022)Liu, Zhang, Gu, Dai, Shan, Song, Li, Chen, Lin, and Sun] can only be used to the last day’s image.
To address the limitations discussed above, we propose a Decoupling Fusion Network called DeFusion to effectively integrate temporal images of the first three days and parental fertility table indicators for IVF-ET pregnancy prediction. The main contributions are summarized as follows:
DeFusion is the first to integrate the first three days of embryonic development temporal images and parental fertility table indicators for pregnancy prediction.
We propose a spatial-temporal position encoding for fusing temporal embryo images. Moreover, we apply a table transformer to extract tabular information from fertility indicators.
We propose a novel decoupling fusion network to fuse multi-modal information more finely grained by decoupling information from different modalities into modality related and unrelated feature.
2 Method
In the field of medical multi-modal fusion, the final fusion features are obtained by complementing the unique features and enhancing the common features of different modalities. However, there is a complex relationship between the features of the same modality and different modalities, which is not a simple linear relationship. So it is difficult to be captured by the model. Inspired by the decoupling operation in [Dong et al.(2023)Dong, Nejjar, Sun, Chatzi, and Fink] [Li et al.(2023a)Li, Fei, Liao, Zhao, Teng, seng Chua, Ji, and Li], we use the decoupling fusion strategy explicitly decouples the features of different modalities into unique and common features, transforming modal features from disorder to order, simplifying the relationships between features and better modeling the complex interactions between modalities. So we propose the decoupling fusion module, a simple and effective multi-modal fusion module as showed in Fig. 2. The input information of the model are temporal grayscale embryo images and fertility table indicator. Embryo images are denoted as , where denote different days, and and denote the height and the width of an image, respectively. Table indicators are denoted as , where denotes the number of indicators.
2.1 Image Extractor
To fuse the first three days embryo images for pregnancy prediction, we design a temporal image fusion network (image extractor). This network consists of three parts as shown in Fig. 3: a backbone aiming at extracting image features; a spatial-temporal position encoding (STPE) to obtain spatial information within a single image and temporal information among different images; and a Transformer [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] that combine STPE to fuse temporal image features. Firstly, we use the backbone to extract embryo image features from the first three days. Then, we reshape the image features into tokens and add positional information to them. We encode image features as spatial position encoding through a convolution layer: . By using convolution operation to capture local spatial features, we can obtain spatial position information within a single image [Chu et al.(2021)Chu, Tian, Zhang, Wang, and Shen]. Image features are encoded as temporal position encoding through a pooling layer and a series of recurrent layers: . By using recurrent operations to capture temporal dependencies, we can obtain temporal position information among different images [Hochreiter and Schmidhuber(1997)]. To align image features, the needs to be replicated to get . The and are integrated through a position encoding attention:
where . The final is as follows:
where . After obtaining the STPE, integrating the tokens obtained through image feature reshaping into the transformer encoder can effectively fuse temporal embryo images for pregnancy prediction.
2.2 Table Extractor
Inspired by TabTransformer [Huang et al.(2020)Huang, Khetan, Cvitkovic, and Karnin] (table extractor), we extract table information of fertility table indicators by transformer [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin]. In order to adapt the table information to the transformer, we specify tabular embedding as a linear layer to upscale the table features to . Next, we construct a series of transformer layer with multi-head self-attention (MHSA) to extract table information:
here, means layer-norm and is the linear layer.
2.3 Decoupling Fusion Module
We use an image feature extractor and a table feature extractor to extract temporal image features and table features. And we denote their outputs as and , respectively. In order to fuse information from different modalities at a finer granularity, as shown in Fig. 2, we propose a decoupling fusion module that decouples the feature of different modalities into related (common) feature and unrelated (unique) feature. We extract related feature and between modalities through shared , and extract unrelated feature and through and . We can decouple the features of different modalities by using cross reconstruction method [Ji et al.(2021)Ji, Zhou, Wang, Wu, Loy, Cao, and Xu, Liu et al.(2023b)Liu, Wei, Li, and Zhou]. The cross reconstruction loss is as follows:
where is the L1-norm, is the dimensionality of the feature. We obtain the final pregnancy prediction result by fusing the decoupled common and unique features: . Finally, we apply cross entropy loss to minimize the difference between the predicted results and the true labels. The classification loss is as follows:
where represent true labels, is the size of a batch. The overall loss function is as follows:
3 Experiments
3.1 Dataset
The first dataset used in the research is from Southern Medical University, with a total of 4046 valid embryo transfer cases. Each case includes both image data and tabular data. The image data are the first three days’ microscopic images of embryo development. The tabular data include 22 parental fertility indicators. The label of each example is positive or negative, representing whether having fetal hearts successfully or not. We conduct a 5-fold cross validation on the dataset.
The second dataset comes from the Peking University International Competition on Ocular Disease Intelligent Recognition (ODIR) [ODIR(2019)], and the original task was to classify eye diseases through image uni-modality. In order to make the dataset applicable to multiple modalities, we extract a total of 3500 cases of image modality and table modality information for eye disease prediction. Among them, the image modality consists of a eye image, and the table modality consists of 8 indicators converted from keywords. We have already made this dataset accessible to the public as a new multi-modal dataset. We conduct a 4-fold cross validation on the dataset.
3.2 Evaluation Metric and Experimental Settings
In the experiments, we evaluate the performance with Accuracy, Area Under the ROC (AUC), and F1-score. AUC is a comprehensive metric to evaluate prediction accuracy. F1-score is an index taking into account the precision and recall of the model predictions.
We implement our method with PyTorch on a Nvidia GeForce RTX 2080ti graphics processing unit (GPU). In addition, since the image class tokens and the table class tokens need to share an encoder, we align them with a linear layer. The learning rate of the image extractor is 1e-6, while the learning rate of the table extractor model is 1e-4. The learning rate of the DeFusion model is 1e-5. The above models all use the Adam [Kingma and Ba(2014)] optimizer. As in Section 3, , , , and .
| Modality | Method | AUC | F1 | Accuracy |
| Table | MLP | 0.684(0.006) | 0.649(0.012) | 0.641(0.013) |
| SVM | 0.690(0.006) | 0.643(0.011) | 0.634(0.011) | |
| Adaboost | 0.708(0.009) | 0.631(0.016) | 0.640(0.015) | |
| TabNet | 0.702(0.014) | 0.643(0.009) | 0.634(0.010) | |
| Ours(TabTransformer) | 0.713(0.012) | 0.661(0.011) | 0.653(0.006) | |
| Image | ResNet+Add | 0.554(0.008) | 0.584(0.012) | 0.592(0.014) |
| ResNet+LSTM | 0.572(0.014) | 0.591(0.018) | 0.600(0.026) | |
| ResNet+Learnable | 0.589(0.012) | 0.597(0.013) | 0.598(0.020) | |
| ResNet+SinCos | 0.602(0.011) | 0.607(0.014) | 0.612(0.021) | |
| Ours(ResNet+STPE) | 0.617(0.012) | 0.621(0.022) | 0.631(0.016) | |
| w/o SPE | 0.614(0.013) | 0.611(0.013) | 0.627(0.015) | |
| w/o TPE | 0.596(0.011) | 0.607(0.008) | 0.613(0.007) | |
| w/o PEAttention | 0.600(0.009) | 0.605(0.007) | 0.613(0.020) | |
| w/o STPE | 0.584(0.016) | 0.603(0.015) | 0.619(0.021) | |
| with ViT | 0.604(0.009) | 0.609(0.013) | 0.613(0.016) | |
| Image and Table | MMBE | 0.723(0.004) | 0.682(0.021) | 0.681(0.011) |
| MOAB | 0.719(0.003) | 0.681(0.009) | 0.674(0.011) | |
| SFusion | 0.718(0.005) | 0.667(0.013) | 0.658(0.013) | |
| ConGraph | 0.718(0.011) | 0.649(0.009) | 0.640(0.009) | |
| HMCAT | 0.723(0.011) | 0.655(0.014) | 0.646(0.012) | |
| Ours(DeFusion) | 0.746(0.003) | 0.689(0.017) | 0.691(0.010) | |
| w/o Decoupling Module | 0.715(0.007) | 0.689(0.013) | 0.681(0.014) | |
| with TabNet | 0.735(0.010) | 0.681(0.011) | 0.683(0.013) |
3.3 Baseline Methods
We compare our method with baseline methods as shown in Table 1, our model achieves superior performance in all the evaluation metrics. Firstly, for table modality, we compare a SVM [Dehghan et al.(2024)Dehghan, Rabiei, Choobineh, Maghooli, Nazari, and Vahidi-Asl] and a Adaboost [Dehghan et al.(2024)Dehghan, Rabiei, Choobineh, Maghooli, Nazari, and Vahidi-Asl] as uni-modal models based on parental tabular fertility indicators. We also compare TabNet [Arik and Pfister(2021)], which is a neural network model specifically designed for tabular classification tasks. The TabTransformer method we use has the best performance. Secondly, for image modality, by comparing temporal image fusion models based on Add, LSTM [Hochreiter and Schmidhuber(1997)], and Transformer with different positional encodings (sin-cos and learnable) [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin], our STPE achieve optimal performance in the transformer-based temporal image fusion strategy. Thirdly, for image and table modalities, we compare the recent multi-modal baseline methods have been introduced in Appendix A and B. Our method, as a new category fusion approach, achieves optimal results.
3.4 Ablation Study
We conduct ablation experiments to evaluate the contribution of each module in our model in Table 1. We first ablate the Spatial-Temporal Position Encoding (STPE) including Spatial Position Encoding (SPE), Temporal Position Encoding (TPE) Position Encoding Attention (PEAttention) titled as w/o SPE, w/o TPE, w/o PEAttention and w/o STPE, respectively. In addition, by comparing the performance of ResNet [He et al.(2016)He, Zhang, Ren, and Sun] and ViT (with ViT) [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby], we choose ResNet as the backbone of the image. Similarly, we evaluate the contribution of the Decoupling Module through w/o Decoupling Module. In addition, by comparing TabNet [Arik and Pfister(2021)] (with TabNet) and TabTransformer, we choose TabTransformer as the table extractor.
3.5 Generalization
Although the proposed DeFusion model is designed for pregnancy prediction, the principles behind it are universal and can be transferred to other multi-modal medical image analysis tasks. We extend DeFusion for multi-modal eye disease prediction on the ODIR dataset. The final prediction results are shown in the Table 2. Although our model doesn’t achieve the highest accuracy, it performs best in the AUC metric, indicating that our model performs better in terms of overall performance.
| Method | AUC | F1 | Accuracy |
| MMBE | 0.836(0.007) | 0.765(0.009) | 0.766(0.007) |
| MOAB | 0.827(0.004) | 0.756(0.005) | 0.751(0.004) |
| SFusion | 0.791(0.030) | 0.755(0.013) | 0.755(0.010) |
| ConGraph | 0.836(0.009) | 0.756(0.008) | 0.751(0.008) |
| HMCAT | 0.835(0.006) | 0.773(0.017) | 0.772(0.010) |
| Ours(DeFusion) | 0.842(0.004) | 0.772(0.009) | 0.770(0.009) |
| w/o Decoupling | 0.825(0.003) | 0.763(0.014) | 0.759(0.014) |
3.6 Visualization
As Fig. 4 shows, we output t-SNE [van der Maaten and Hinton(2008)] results and average Pearson correlation coefficient (PCC) matrix [Sverko et al.(2022)Sverko, Vrankic, Vlahinic, and Rogelj] of (img_related), (tab_related), (img_unrelated) and (tab_unrelated) from the decoupling test set. The PCC is between 0 and 1, with a larger value indicating greater relevance. The overlap between the points of and after t-SNE dimensionality reduction and the high PCC value between and indicate that the model successfully capture relevant and overlapping information between the two modalities. On the contrary, and are well separated, indicating it capture the information that is independent and complementary between the two modalities. These prove the effectiveness of the decoupling module.
4 Conclusion
This paper proposes a Decoupling Fusion Network called DeFusion to integrate the multi-modal information of temporal embryo images and parental fertility table indicators for IVF-ET pregnancy prediction. The superior performance suggest that our model can provide valuable assistance for the selection of embryos for transplantation. And the effectiveness of the decoupling fusion module has been demonstrated through visualization and generalization experiments. In the future, we will optimize the decoupling module and expand it to more datasets.
Appendix A Related Work
Fusion of heterogeneous information from multi-modal data can effectively enhance model performance, which is a key project in the medical field of multi-modal learning [Cui et al.(2023)Cui, Yang, Wang, Zhao, Asad, Coburn, Wilson, Landman, and Huo]. Decision-level fusion and feature-level fusion are two main strategies for multi-modal fusion. The decision-level fusion employs averaged, weighted voting or majority voting [Holste et al.(2021)Holste, Partridge, Rahbar, Biswas, Lee, and Alessio] to integrate the outputs of uni-modal models, so as to make the final multi-modal output.
Although decision-level fusion is simple to implement, it cannot capture the interactions between hidden features from different modalities. The feature-level fusion fuses the heterogeneous multi-modal data by projecting extracted features into a compact and information-rich multi-modal hidden representation space. Feature-level fusion mainly includes simple-operation based, tensor-based, transformer-based and graph-based methods. The simple-operation method performs concatenation, element addition and element multiplication operations. [Zhou et al.(2021)Zhou, Zhang, Zhu, Lan, Fu, Wang, and Wen] use two branch encoders to extract image and non-image information, and fuse the extracted information on key point through simple operation for COVID-19 patient severity prediction. The tensor-based method performs outer product between multi-modal feature vectors to form higher-order co-occurrence matrices, which provide more informative information than these features alone. [Zolotarev et al.(2024)Zolotarev, Khan, Slabaugh, and Roney](MOAB) use a deep orthogonal fusion model to predict the atrial fibrillation from different multi-modal data. The attention mechanism in the transformer-based method has the ability to aggregate features in different feature spaces, making it very suitable for multi-modal alignment and fusion. [Liu et al.(2023c)Liu, Wei, Li, and Zhou](SFusion) apply transformer to fuse different modalities of brain imaging for tumor segmentation; [Li et al.(2023b)Li, Jiang, Lu, Li, and Xia](HMCAT) use a model based on the cross attention transformer that integrates pathological and radiological images for cancer prediction. Graph-based modeling and inference can provide a deeper understanding of disease information by discovering complex relationships between hidden disease tissue regions. [Ding et al.(2024)Ding, Hsu, and Liu](ConGraph) transform images and non-images into graph nodes based on fully connected graph attention network, and fuse information among nodes to predict Pakinson’s disease.
Appendix B Baseline Methods
B.1 Baseline methods with image modality
ReNet+Add applys ResNet to extract image information from the first three days of embryonic development, where . Then, , and are fused through an addition operator. Finally, we use a classifier consisting of three non-linear layers for pregnancy prediction. ReNet+LSTM replaces the addition operator with LSTM on the basis of ReNet+Add. No Position, Learnable and Sin-Cos are the results of replacing our proposed spatial-temporal position encoding with different position encoding.
B.2 Baseline methods with both image modality and table modality
As shown in Fig. A.1, we compare different multi-modal fusion methods. To ensure fairness in comparison, all methods use the same backbone. In addition, according to the structural characteristics of different models, TransformerFusion (SFusion [Liu et al.(2023c)Liu, Wei, Li, and Zhou], HMCAT [Li et al.(2023b)Li, Jiang, Lu, Li, and Xia]) and GraphFusion (ConGraph [Ding et al.(2024)Ding, Hsu, and Liu]) use all image and table tokens, while AddFusion (MMBE [Liu et al.(2022)Liu, Zhang, Gu, Dai, Shan, Song, Li, Chen, Lin, and Sun]), TensorFusion (MOAB [Zolotarev et al.(2024)Zolotarev, Khan, Slabaugh, and Roney]), and our DecouplingFusion (DeFusion) only use image class token and table class token.
As shown in Fig. A.1 (a) and (b), AddFusion and TensorFusion directly fuse the class tokens of the two modalities using an addition operation and an outer product operation, respectively. As shown in Fig. A.1 (c) and (d), TransformerFusion and GraphFusion use the Graph Attention Network [Velickovic et al.(2017)Velickovic, Cucurull, Casanova, Romero, Lio, Bengio, et al.] and the Transformer [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] as the fusion network to fuse all tokens of the two modalities, respectively. Our method is an innovative multi-modal fusion category.
Appendix C Dataset Samples
In the Fig.A.2 and Table A.1, we present some examples of image modality and table modality used for pregnancy prediction, respectively. The image is obtained by cropping the excess blank space in the original image. We replace the missing value in the table data with the average value and normalize the tabular features with min-max before inputting them into the model.
| Number | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 |
| Female age | 39 | 25 | 44 | 34 | 31 | 36 | 33 | 29 | 28 | 29 |
| Man age | 46 | 24 | 45 | 36 | 35 | 38 | 38 | 32 | 35 | 31 |
| Female BMI | 22.8 | 18.7 | 22.1 | 21.5 | 18.5 | 23.9 | 23.4 | 18 | 25 | 20.4 |
| AFC | 10 | - | - | 8 | 18 | 9 | 12 | 16 | 10 | 8 |
| Number of obtained oocytes | 5 | 22 | 3 | 10 | 17 | 14 | 5 | 11 | 7 | 7 |
| Number of mature oocytes | 4 | 14 | 3 | 10 | 15 | 10 | 5 | 8 | 7 | 7 |
| Available embryos | 3 | 2 | 3 | 7 | 7 | 3 | 4 | 3 | 5 | 7 |
| High-quality embryos | 2 | 1 | 1 | 7 | 5 | 1 | 2 | 2 | 4 | 4 |
| HCG Day E2 | 1037 | 2271 | 1782 | 1483 | 3118 | 3693 | 731.9 | 1891 | 1309 | 3767 |
| HCG intimal thickness | 7.5 | 15.5 | 8 | 9.3 | 9.4 | 11.7 | 9.8 | 15 | 17.5 | 12.5 |
| FSH | 8 | - | - | 7.81 | 7.22 | 4.07 | 4.61 | 6.53 | 7.3 | - |
| Infertility years | 2 | - | - | 9 | - | 7 | 10 | 1 | 3 | 4 |
| Volume BT | 1.4 | 4.1 | 4 | 2.6 | 2 | 1.5 | 1.5 | 1 | 2.5 | 0.3 |
| Concentration BT | 6 | 2 | 25 | 20 | 0.01 | 3 | 4 | 10 | 18 | 40 |
| Non forward movement BT | 5 | 10 | 5 | 15 | 0.5 | 5 | 3 | 10 | 10 | 5 |
| Inactivity BT | 90 | 80 | 85 | 60 | 0.5 | 90 | 95 | 85 | 80 | 75 |
| Forward movement BT | 5 | 10 | 10 | 25 | 0.5 | 5 | 2 | 5 | 10 | 10 |
| Volume AT | 0.1 | 0.1 | 0.6 | 0.2 | 0.2 | 0.15 | 0.2 | 0.15 | 0.3 | 0.3 |
| Concentration AT | 1 | 2 | 3 | 8 | 0.01 | 1 | 1 | 1 | 1 | 3 |
| Non forward movement AT | 20 | 10 | 5 | 5 | 0.5 | 10 | 20 | 20 | 20 | 10 |
| Inactivity AT | 30 | 25 | 5 | 5 | 0.5 | 10 | 10 | 40 | 10 | 5 |
| Forward movement AT | 50 | 65 | 90 | 90 | 0.5 | 80 | 70 | 30 | 60 | 85 |
Appendix D Supplementary Experiments
D.1 Table Modality Experiments
As shown in Fig.A.3, it is a detail of the table extractor (TabTransformer).
D.2 Image Modality Experiments
Table A.2 shows the performance with image modality only. We use images of embryo development on the first, second, and third day for pregnancy prediction. And the backbone is ResNet. The results indicate that as the embryo continues to develop, it provides greater assistance in predicting pregnancy.
| Modality | Method | AUC | F1 | Accuracy |
| Image | ResNet(First Day) | 0.476(0.004) | 0.515(0.009) | 0.502(0.009) |
| ResNet(Second Day) | 0.565(0.007) | 0.572(0.013) | 0.560(0.013) | |
| ResNet(Third Day) | 0.593(0.007) | 0.598(0.012) | 0.595(0.010) | |
| Ours(ResNet+STPE)(Three Days) | 0.617(0.012) | 0.621(0.022) | 0.631(0.016) |
D.3 Computational Complexity
As shown in the Table A.3, we compare the computational complexity of multi-modal fusion methods. Although our model is not optimal in terms of computational complexity, compared to other methods, our approach still achieves competitive results in terms of computational complexity while achieving optimal accuracy.
| Method | MMBE | MOAB | SFusion | ConGraph | HMCAT | Ours(DeDusion) |
| GFLOPs(↓) | 12.66 | 13.30 | 13.66 | 14.01 | 16.19 | 12.67 |
D.4 Uni-modal Experiment on The ODIR
As shown in the Table A.4, we compare the uni-modal method on the ODIR dataset for eye disease prediction.
| Modality | Method | AUC | F1 | Accuracy |
| Image | ResNet | 0.714(0.016) | 0.677(0.015) | 0.674(0.015) |
| ViT | 0.736(0.006) | 0.685(0.013) | 0.680(0.014) | |
| Table | SVM | 0.802(0.003) | 0.722(0.001) | 0.714(0.007) |
| TabNet | 0.793(0.024) | 0.708(0.023) | 0.701(0.015) | |
| MLP | 0.793(0.001) | 0.731(0.007) | 0.727(0.003) |
Appendix E Interpretability
In order to analyze the reasons for the success of the model, we conduct interpretability analysis on the model, mainly manifest in two aspects. Firstly, as shown in Fig. A.4, we use SHAP [Lundberg and Lee(2017)] to output the importance ranking of clinical indicators in our model, which focuses more on features such as female age, high-quality embryos and so on. Among them, A.4(a) ranks the importance of table features. A.4(b) is a beeswarm, which depicts the SHAP values of each sample under different features. Secondly, we use the Grad-Cam [Selvaraju et al.(2017)Selvaraju, Cogswell, Das, Vedantam, Parikh, and Batra] to visualize the class activation maps of the first three days of embryonic development images in Fig. A.5. We can see the areas that our model focuses on are the edges of embryonic cells. These interpretable results are consistent with the experience of obstetricians and gynecologists.

### v2

@@
 In order to perform the embryos morphological grading task, [Wu et al.(2021)Wu, Yan, Li, Li, Wang, Chang, Yu, Jin, Ma, Luo, et al.] and [Liu et al.(2023a)Liu, Li, Dai, Shan, Zhang, Zhuang, Lee, Wong, Yue, Huang, et al.] apply convolutional neural networks (CNN) based on static embryo images; [Wang et al.(2024)Wang, Wang, Gao, Chen, Gao, Ma, Jiang, Yang, Feng, Zhang, et al.] and [Lukyanenko et al.(2021)Lukyanenko, Jang, Wei, Struyven, Kim, Leahy, Yang, Rush, Ben-Yosef, Needleman, et al.] apply transformer and two-stream neural network based on time-lapse microscopy (TLM) images, respectively.
-In addition, [Cheng et al.(2024)Cheng, Chen, Hu, Meng, Liu, Chen, Wu, and Ying] fuse multi-focal images to predict outcome of blastocyst.
+In addition, [Cheng et al.(2024)Cheng, Chen, Hu, Meng, Liu, Chen, Wu, and Ying] fuse multi-focal images to predict grade of blastocyst.
 The performance of these methods outperform laboratory technicians, because embryos grading is completely based on morphological information and the salient morphological characteristics among different grades are easily distinguishable for machines.
@@
 So it is difficult to be captured by the model.
-Inspired by the decoupling operation in [Dong et al.(2023)Dong, Nejjar, Sun, Chatzi, and Fink] [Li et al.(2023a)Li, Fei, Liao, Zhao, Teng, seng Chua, Ji, and Li], we use the decoupling fusion strategy explicitly decouples the features of different modalities into unique and common features, transforming modal features from disorder to order, simplifying the relationships between features and better modeling the complex interactions between modalities.
+Inspired by the decoupling operation in [Dong et al.(2023)Dong, Nejjar, Sun, Chatzi, and Fink] [Li et al.(2023a)Li, Fei, Liao, Zhao, Teng, seng Chua, Ji, and Li], we use the decoupling fusion strategy explicitly decouples the features of different modalities into unique and common features, which is a shift from entangled features to disentangled ones, simplifying the relationships between features and better modeling the complex interactions between modalities.
 So we propose the decoupling fusion module, a simple and effective multi-modal fusion module as showed in Fig.
@@
 3: a backbone aiming at extracting image features; a spatial-temporal position encoding (STPE) to obtain spatial information within a single image and temporal information among different images; and a Transformer [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] that combine STPE to fuse temporal image features.
-Firstly, we use the backbone to extract embryo image features from the first three days.
+Firstly, we use the backbone to extract embryo image features from the first three days, where is the number of channels and is the scaling factor.
 Then, we reshape the image features into tokens and add positional information to them.
@@
 The cross reconstruction loss is as follows:
-where is the L1-norm, is the dimensionality of the feature.
-We obtain the final pregnancy prediction result by fusing the decoupled common and unique features: .
+where is the L1-norm, is the dimensionality of the feature, and are decoder.
+We obtain the final pregnancy prediction result by fusing the decoupled common and unique features: , which is composed of three layers of MLP.
 Finally, we apply cross entropy loss to minimize the difference between the predicted results and the true labels.
@@
 where represent true labels, is the size of a batch. The overall loss function is as follows:
+where is a hyperparameter in the loss function.
 3 Experiments
@@
 In the Fig.A.2 and Table A.1, we present some examples of image modality and table modality used for pregnancy prediction, respectively.
-The image is obtained by cropping the excess blank space in the original image.
-We replace the missing value in the table data with the average value and normalize the tabular features with min-max before inputting them into the model.
+When processing image data, we use the following data enhancement during training and testing: resize to 256 pixels and then center cropping to 224 pixels.
+Finally, normalization with a mean of 0.566 and a variance of 0.063 is used, and the mean and variance were obtained by statistics of the whole data set.
+When processing tabular data, we use the average of features to replace missing features.
+Then we input the features into the neural network after min-max normalization.
 | Number | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 |
@@
 | GFLOPs(↓) | 12.66 | 13.30 | 13.66 | 14.01 | 16.19 | 12.67 |
+| Training Time(↓) | 50 | 91 | 65 | 68 | 79 | 64 |
+| Inference FPS(↑) | 5.05 | 2.69 | 3.34 | 3.54 | 2.79 | 3.51 |
 D.4 Uni-modal Experiment on The ODIR
@@
 | MLP | 0.793(0.001) | 0.731(0.007) | 0.727(0.003) |
+D.5 Generalization of Other Dataset
+We collect a dataset of 218 cases from Guangzhou Women and Children’s Medical Center as an additional test set to test our DeFusion model, with 56 pregnant cases and 162 non-pregnant cases in this dataset.
+As shown in the Table A.5, our model has certain generalization ability without fine tuning.
+| Modality | Method | AUC | F1 | Accuracy |
+| Image+Table | Ours(DeFusion) | 0.616 | 0.660 | 0.642 |
 Appendix E Interpretability
@@
 These interpretable results are consistent with the experience of obstetricians and gynecologists.
+Appendix F Limitation
+Although our study is superior to other methods, there are limitations to our method and data set.
+For our method, our premise is that there is correlation between the multi-modal data.
+If there is no correlation between the multi-modal data, our method may not work well.
+In addition, when decoupling unique features and common features, we only use a simple cross-reconstruction loss constraint, which is very weak.
+Although the decoupling visualization by t-SNE and Pearson correlation coefficient proves the effectiveness of the decoupling method, adding stronger loss function constraints to the decoupling process may make the decoupling process smoother.
+Finally, because we add the decoupling module to the multi-modal fusion process, our computational complexity is higher than the simple ADD fusion, which is not conducive to our deployment of the model to the end-to-end device.
+For our dataset, we only collected 4046 cases of data, which is not enough for deep learning.
+And, there are some missing values in our clinical indicator data, which is unfavorable to the prediction of results.
+In addition, our image data is three images taken every other day using a normal microscope, and many studies now use time-lapse microscopes, which can acquire images at the hour or even minute level, so our image data lacks a lot of temporal information compared to other studies.